OpenStudy (anonymous):

x^2 + 4x +1 = 0 Solve using quadratic formula Medals awarded :)

OpenStudy (amoodarya):

x^2+4x+1= x^2+4x+4-4+1 =(x+2)^2-4+1

OpenStudy (amoodarya):

=(x+2)^2-3=0 so x+2=+-sqrt3

OpenStudy (anonymous):

Sorry I do not understand that.

OpenStudy (anonymous):

Have done and i dont seem to get the right answer

OpenStudy (anonymous):

-0.27 and -3.73 are the correct answers

OpenStudy (amoodarya):

x1 = -0.2679 x2 = -3.7321

OpenStudy (anonymous):

but i dont know how to get to them.

OpenStudy (amoodarya):

sqrt3=1.73 so x=-2+-1.73 x=-.27 x=-3.73

OpenStudy (anonymous):

i need the whole method from substituting numbers into equation of x=-b etc

OpenStudy (nurali):

\[\frac{ -b+-\sqrt b^2-4ac }{ 2a }\] \[\frac{ -4+-\sqrt (4)^2-4*1*1 }{ 2*1 }\] \[\frac{ -4+-\sqrt 12 }{ 2 }\] \[x1=\frac{ -4+\sqrt12 }{ 2 }, x2=\frac{ -4-\sqrt12 }{ 2 }\]

OpenStudy (anonymous):

Thanks! :)

OpenStudy (nurali):

My Pleasure.